# Constraints from Entanglement Wedge Nesting for Holography at a Finite Cutoff

### v1

@@ Abstract
We explore constraints that arise from associating an entanglement wedge (EW) to subregions of a cutoff boundary at a finite distance in AdS/CFT, using a subcritical end-of-the-world (ETW) brane acting as the cutoff.
@@ 1.1 Motivation and Background
One attempt at generalizing the duality between quantum gravity in a bulk AdS spacetime and a quantum field theory on the asymptotic boundary is to consider spacetimes with boundaries at a finite proper distance. Understanding holography in such a setting would enable to study quantum gravity in finite local subregions, but also to potentially embed cosmologies in holography Gubser:1999vj ; Cooper:2018cmb . More generally, such ideas can be understood in the framework of the surface/state correspondence Miyaji:2015yva ; Miyaji:2015fia , which assigns a quantum state |$\ket{\phi(\Sigma)}$ to convex codimension-two submanifolds $\Sigma$ in a gravitational spacetime.
Concrete realizations for the surface/state correspondence can be found, apart from tensor networks which motivated the correspondence, for example in the study of $T\bar{T}$-deformations of holographic CFTs Zamolodchikov:2004ce ; Jiang2019LecturesOS ; Bonelli:2018kik ; Dubovsky:2017cnj ; Grieninger:2019zts ; Gross:2019ach ; Cavaglia:2016oda . It has been proposed that a $T\bar{T}$ deformation of a holographic CFT should be understood as “moving the boundary into the bulk” McGough:2016lol ; Guica:2019nzm ; Lewkowycz:2019xse . Furthermore, the appearance of cutoff surfaces have also played a key role in double holography Takayanagi:2011zk ; Almheiri:2019hni ; Chen:2020uac ; Chen:2020hmv ; Hernandez:2020nem ; Grimaldi:2022suv ; Neuenfeld2021DoubleHA ; Neuenfeld:2021wbl ; Omiya:2021olc ; Geng:2023iqd , where a convex cutoff surface in the form of an end-of-the-world (ETW) brane appears. Doubly-holographic models with AdS branes start by introducing the dual of a holographic $d$-dimensional boundary BCFT in terms of a $d+1$-dimensional bulk AdS spacetime. [^1] BCFTs are conformal field theories living on a space with boundary at which conformal boundary conditions are imposed. However, at least for a BCFT2 one needs to introduce boundary degrees of freedom which break conformal invariance in order to obtain interesting models, e.g. by coupling an SYK model to the BCFT boundary. We will will thus use the term BCFT more loosely to also account for situations like this. The boundary of the BCFT is continued into the bulk by an ETW brane Takayanagi:2011zk which removes part of the naive bulk spacetime dual to a CFT without boundary.
@@
In addition, in standard AdS/CFT we have a notion of subregion/subregion duality Headrick:2014cta ; Akers:2016ugt ; Akers:2017ttv ; Chen:2021lnq ; Leutheusser:2022bgi . More precisely, a boundary subregion $U$ is dual to the entanglement wedge $\mathcal{W}_{E}(U)$ which is defined as the causal past and future of the partial bulk Cauchy slice bounded by $U$ and the corresponding RT surface $\chi(U)$. Moreover, given a subregion $U^{\prime}\subset U$, $U^{\prime}$ is dual to $\mathcal{W}_{E}(U^{\prime})$ which must be a subset of $\mathcal{W}_{E}(U)$. This follows from the fact that tracing out parts of $U$ to $U^{\prime}$ acts as a quantum channel on the reduced density matrix $\rho_{U}$. This property of entanglement wedges is called entanglement wedge nesting (EWN).
The purpose of this paper is to first carefully describe the subtleties in defining entanglement wedges for cutoff subregions $B$, and second, to the non-trivial conditions on the relative location of $B$ with respect to another subregion $A$ (disjoint from $B$) imposed by EWN. For concreteness, we will focus on the simplest geometry which allows for non-trivial statements, namely AdS3 and Planar BTZ cut off by a hyperbolic ETW brane. Throughout this paper, we will take the perspective that one can freely choose a subregions $A$ and $B$ in much the same way as one can choose a subregion on a standard holographic boundary. [^3] The reader should be cautious to not confuse this with the closely related notion of islands. Although the case of finding islands occurs in the connected phase the issue of knowing where on the brane the extremal surface is anchored is dynamically determined as part of the extremization problem. In our case we are considering fixing a brane region and fixing a boundary region independently. In a certain sense we are considering Dirichlet boundary conditions on the brane whereas island setups use Neumann conditions. However, the theory on the cutoff surface might in general be gravitational as is the case in doubly holography Karch:2000ct . This naively poses the question of how to define cutoff subregions in a diffeomorphism invariant way. In this paper, we restrict ourselves to the strict $N\to\infty$ limit, such that gravity is effectively turned off in the bulk. In this case subregions can be defined uniquely. [^4] The authors of Bousso:2022hlz proposed the notion of generalized entanglement wedges for gravitational theories. Their proposal only works for subregions of a time-reflection symmetric Cauchy slice and thus is not applicable to our configurations. Extending this discussion to $1/N$ corrections, while interesting, will not be pursued here.
@@
We will assume that the RT formula in its simplest formulation, Eq. (1), holds. Given this starting point, the RT surface which computes $S(A\cup B)$ for two disconnected, spacelike subregions $A$ and $B$ consists of two disconnected pieces which can be in one of two phases. Either, each connected piece can be homologous to one of $A$ and $B$. In this case, which we will call the disconnected phase, the mutual information $S(A)+S(B)-S(A\cup B)$ vanishes. The second phase has the RT surface connect the subregions $A$ and $B$ to each other. Consequently, the mutual information does not vanish. We will call this the connected phase. We will then construct the entanglement wedges $\mathcal{W}_{E}$ for the regions $A$, $B$, and $A\cup B$. As we will argue below, while the definitions of $\mathcal{W}_{E}(A)$ and $\mathcal{W}_{E}(B)$ are straight forward and agree with the intersection of entanglement wedge in the full geometry cut off by the brane, the definition of the connected entanglement wedge is more subtle and requires us to be careful about the location and direction at which the RT surface ends on the cutoff surface.
@@ 2.1 AdS3 Poincare Patch with an ETW Brane
where $L$ is the AdS radius, $z>0$ is the radial bulk coordinate and $z=0$ is the conformal boundary which is parameterized by $t,x\in\mathbb{R}$.
The ETW brane is a co-dimension one surface in the bulk which satisfies the the metric equation of motion of the Lorentzian action [^5] Note in our setup in Poincare AdS the boundary of the bulk cutoff by the brane consists of the conformal boundary as well as the ETW brane so we should also include a Gibbons-Hawking term of the form $\int_{\text{bdry}}d^{d}y\sqrt{-h_{\text{bdry}}}K_{\text{bdry}}$. For simplicity we omit explicitly writing it this in the action and focus on the brane and bulk terms. Takayanagi:2011zk
$K_{ab}-Kh_{ab}+Th_{ab}=0,$ (4)
$K_{ab}=\frac{T}{d-1}h_{ab}.$ (5)
It is straightforward to check that in Poincare coordinates the solution for an ETW brane that intersects the conformal boundary at $x=0$ is given by
@@
where $\theta_{0}\in(0,\pi/2]$ is a parameter that is related to the brane tension via $TL=\sin\theta_{0}$. When the tension is zero the ETW brane cuts off half of the AdS space and when $T=L^{-1}$ the ETW brane coincides with the conformal boundary. The limit of $T\lesssim\frac{1}{L}$, i.e., the limit in which the brane approaches the “would be conformal boundary” is referred to as the critical limit. The critical limit is physically relevant in the discussion of gravity induced on the brane, particularly in higher dimensions. It is a limit in which bulk gravitons become localized on the brane and one can consider a conventional lower dimensional gravity theory on the world volume of the brane Karch:2000ct ; Geng:2020qvw ; Neuenfeld:2021wbl . In the three-dimensional case discussed here, a theory of dilaton gravity gets induced on the brane, and the critical limit is the limit in which the cutoff scale of the brane theory becomes small Neuenfeld:2023svs .
@@ 2.3 General Classification of Extremal Surfaces in Poincare AdS3
Nonetheless, for $|k|$ above a certain threshold, the extremal surface becomes timelike. Here, we consider only spacelike separated points with respect to the bulk metric, as expected in computations of entanglement entropy for spacelike separated subregions. [^7] For work on timelike entanglement entropy see Doi:2022iyj ; Liu:2022ugc ; Doi:2023zaf ; Li:2022tsv . In this case we have that $|\Delta t|<\sqrt{\Delta x^{2}+\Delta z^{2}}$, where $\Delta x$ and $\Delta z$ are the coordinate distances between the RT endpoints. This places an upper bound on the parameter $|k|<k_{c}$, with $k_{c}$ given by
@@ 3.1 Condition for EWN
where the right hand side defines the additional bounds $\Delta t_{i,EWN}$.
@@ 4 Entanglement Wedge Nesting in a Planar BTZ Black Hole Geometry
In this section, we will study conditions for entanglement wedge nesting for a two-sided planar BTZ black hole geometry with one exterior having an ETW brane by adapting the formalisms developed in the previous sections in AdS3. BTZ black hole geometry has also been of great interest in a variety of contexts ranging from black hole microstates Hartman:2013qma ; Kourkoulou:2017zaj ; Miyaji:2021ktr , brane world cosmologies Cooper:2018cmb ; Antonini:2019qkt ; Waddell:2022fbn , as well as constraining DGP couplings in brane-world setups Lee:2022efh , which motivates their study using our techniques.
@@ 4.1 Two-Sided BTZ Black Holes with an ETW Brane
We now introduce an ETW brane in our setup. The ETW brane will be a timelike co-dimension one surface, which satisfies the following equations of motion involving the extrinsic curvature and induced metric
@@ 4.4 Conditions for EWN
The tightest interval occurs when we take the limit of the expressions above as $\Xi\to 1$ (i.e. when $a\to\infty$) to give,

### v2

@@ Abstract
We explore constraints that arise from associating an entanglement wedge (EW) to subregions of a cutoff boundary at a finite distance in AdS/CFT, using a subcritical end-of-the-world (ETW) brane acting as a cutoff.
@@ 1.1 Motivation and Background
One attempt at generalizing the duality between quantum gravity in a bulk AdS spacetime and a quantum field theory on the asymptotic boundary is to consider spacetimes with boundaries at a finite proper distance. Understanding holography in such a setting would enable the study of quantum gravity in finite local subregions, and also potentially embed cosmologies in holography Gubser:1999vj ; Cooper:2018cmb . More generally, such ideas can be understood in the framework of the surface/state correspondence Miyaji:2015yva ; Miyaji:2015fia , which assigns a quantum state |$\ket{\phi(\Sigma)}$ to convex codimension two submanifolds $\Sigma$ in a gravitational spacetime.
Concrete realizations for the surface/state correspondence can be found, apart from tensor networks which motivated the correspondence, for example in the study of $T\bar{T}$-deformations of holographic CFTs Zamolodchikov:2004ce ; Jiang2019LecturesOS ; Bonelli:2018kik ; Dubovsky:2017cnj ; Grieninger:2019zts ; Gross:2019ach ; Cavaglia:2016oda ; Pant:2024eno . It has been proposed that a $T\bar{T}$ deformation of a holographic CFT should be understood as “moving the boundary into the bulk” McGough:2016lol ; Guica:2019nzm ; Lewkowycz:2019xse . Furthermore, the appearance of cutoff surfaces have also played a key role in double holography Takayanagi:2011zk ; Almheiri:2019hni ; Geng:2020fxl ; Geng:2020qvw ; Chen:2020uac ; Chen:2020hmv ; Hernandez:2020nem ; Grimaldi:2022suv ; Geng:2021wcq ; Geng:2021iyq ; Neuenfeld2021DoubleHA ; Neuenfeld:2021wbl ; Omiya:2021olc ; Geng:2023iqd ; Geng:2023qwm , where a convex cutoff surface in the form of an end-of-the-world (ETW) brane appears. Doubly-holographic models with AdS branes start by introducing the dual of a holographic $d$-dimensional boundary BCFT in terms of a $d+1$-dimensional bulk AdS spacetime. [^1] BCFTs are conformal field theories living on a space with boundary at which conformal boundary conditions are imposed. However, at least for a BCFT2 one needs to introduce boundary degrees of freedom which break conformal invariance in order to obtain interesting models, e.g. by coupling an SYK model to the BCFT boundary. We will will thus use the term BCFT more loosely to also account for situations like this. The boundary of the BCFT is continued into the bulk by an ETW brane Takayanagi:2011zk which removes part of the naive bulk spacetime dual to a CFT without a boundary.
@@
In addition, in standard AdS/CFT we have a notion of subregion/subregion duality Headrick:2014cta ; Akers:2016ugt ; Akers:2017ttv ; Chen:2021lnq ; Leutheusser:2022bgi . More precisely, a boundary subregion $U$ is dual to the entanglement wedge $\mathcal{W}_{E}(U)$ which is defined as the domain of dependence of the partial bulk Cauchy slice bounded by $U$ and the corresponding RT surface $\chi(U)$. Moreover, given a subregion $U^{\prime}\subset U$, $\mathcal{W}_{E}(U^{\prime})$ is the bulk subregion dual to $U^{\prime}$, and must satisfy $\mathcal{W}_{E}(U^{\prime})\subset\mathcal{W}_{E}(U)$. This follows from the fact that tracing out parts of $U$ to $U^{\prime}$ acts as a quantum channel on the reduced density matrix $\rho_{U}$. This property of entanglement wedges is called entanglement wedge nesting (EWN).
The purpose of this paper is to first carefully describe the subtleties in explicitly defining and constructing entanglement wedges for cutoff subregions $B$, and second, to derive non-trivial conditions on the relative location of $B$ with respect to another subregion $A$ (disjoint from $B$ and living on a conformal boundary) imposed by EWN. For concreteness, we will focus on the simple geometries of AdS3 and Planar BTZ cut off by a hyperbolic ETW brane where computations can be done easily and the constraints we derive can be interpreted in a straightforward manner.
@@
We will assume that the RT formula in its simplest formulation, Eq. (1), holds [^3] The authors of Bousso:2022hlz proposed the notion of generalized entanglement wedges for gravitational theories. Their proposal appears to only work for subregions of a time-reflection symmetric Cauchy slice and thus is not applicable to our configurations. . Given this starting point, the RT surface which computes $S(A\cup B)$ for two disconnected, spacelike subregions $A$ and $B$ consists of two disconnected pieces which can be in one of two phases. Either, each connected piece can be homologous to one of $A$ and $B$. In this case, which we will call the disconnected phase, the mutual information $S(A)+S(B)-S(A\cup B)$ vanishes. The second phase has the RT surface connect the subregions $A$ and $B$ to each other. Consequently, the mutual information does not vanish. We will call this the connected phase. We will then construct the entanglement wedges $\mathcal{W}_{E}$ for the regions $A$, $B$, and $A\cup B$. As we will argue below, while the definitions of $\mathcal{W}_{E}(A)$ and $\mathcal{W}_{E}(B)$ are straight forward and agree with the intersection of entanglement wedge in the full geometry cut off by the brane, the definition of the connected entanglement wedge is more subtle and requires us to be careful about the location and direction at which the RT surface ends on the cutoff surface.
@@
The issue of defining the subregion, $B$, on the cutoff/brane can be subtle depending on the types of boundary condition (BC) we impose on the brane. When discussing the brane/cutoff there appear to be at least two distinct ways in which we might interpret the physics on the brane which depends on whether we choose Dirichlet or Neumann BCs on the brane Geng:2023qwm ; Deng:2023pjs ; Geng:2024xpj . When using Neumann BCs the way the brane is embedded in the higher dimensional AdS geometry is determined dynamically by analyzing equations of motion of the following Lorentzian action [^4] Note in our setup in Poincare AdS the boundary of the bulk cutoff by the brane consists of the conformal boundary as well as the ETW brane so we should also include a Gibbons-Hawking term of the form $\int_{\text{bdry}}d^{d}y\sqrt{-h_{\text{bdry}}}K_{\text{bdry}}$. For simplicity we omit explicitly writing it this in the action and focus on the brane and bulk terms. Takayanagi:2011zk
$K_{ab}-Kh_{ab}+Th_{ab}=0,$ (3)
$K_{ab}=\frac{T}{d-1}h_{ab},$ (4)
which implies the brane/cutoff is a co-dimension one surface in the bulk of constant extrinsic curvature characterized by a tension $T$ and usually one can consider dynamical gravity on the brane as is often done in double holography Takayanagi:2011zk ; Almheiri:2019hni ; Geng:2020fxl ; Geng:2020qvw ; Chen:2020uac ; Chen:2020hmv ; Hernandez:2020nem ; Grimaldi:2022suv ; Neuenfeld2021DoubleHA ; Neuenfeld:2021wbl ; Omiya:2021olc ; Geng:2023iqd ; Geng:2023qwm and brane-world gravity Karch:2000ct ; Geng:2023qwm . This naively poses the questions of how to define cutoff subregions in a diffeomorphism invariant. One possible way to address this by going back to the subregion $A$ which lies on the non-gravitating conformal boundary. Then we consider RT surfaces associated to $A$ which extend to the brane. In the case of Neumann BCs the location the RT surfaces end on the brane is dynamically determined and this in turn can be used to associate a subregion $B$ on the brane given some $A$ on the conformal boundary (i.e. $B$ is not independently chosen from $A$). Since we are considering dynamical gravity on the brane one can change the gravitational dynamics on the brane by adding terms to Eq. (2) (e.g, higher derivative corrections, DGP gravity, etc. on the brane Dvali:2000hr ; Chen:2020hmv ; Chen:2020uac ; Geng:2022tfc ; Geng:2022slq ; Geng:2023qwm ; Lee:2022efh ) these would generally affect where the RT surface ends on the brane and thereby shift the relative location of $A$ and $B$. One might speculate that perhaps one might end up in configurations that violate EWN, in this case the constraints we find may be regarded as providing constraints to gravity on the brane. On the other hand, imposing Dirichlet boundary conditions on the brane one can no longer view the embedding of the brane in AdS as arising from the equation of motion given in Eq. (4). In this case we simply view the brane as a cutoff surface of constant extrinsic curvature where all bulk fields satisfy Dirichlet boundary conditions on the brane, this is more akin to an interpretation of a cutoff in the $T\bar{T}$ context where gravity is effectively turned off on the brane. In such a case we can fix some gauge and naively choose spacelike separated subregions $A$ and $B$ “independently” of each other. In this case what we constrain through EWN is not gravity on the brane as we did in the Neumann case but rather the relative location of $A$ and $B$. The central point here is that regardless of the interpretation we place on the physics of the brane our results simply rely on having two subregions; $A$ on a conformal boundary and $B$ on the brane/cutoff along with a prescription to associate entanglement wedges to these subregions. With this we can derive constraints from EWN and such an investigation on its own will yield interesting results which we summarize in the following section.
@@ 2.1 AdS3 Poincare Patch with an ETW Brane
where $L$ is the AdS radius, $z>0$ is the radial bulk coordinate and $z=0$ is the conformal boundary which is parameterized by $t,x\in\mathbb{R}$. We consider the brane/cutoff as a co-dimension one surface in the bulk which intersects the conformal boundary at $x=0$ and is given by
@@
where $\theta_{0}\in(0,\pi/2]$. As we already discussed in the introduction, when one imposes Neumann boundary conditions on the brane, the brane is a surface which will satisfy Eq. (4) with a tension given by $TL=\sin\theta_{0}$. When the tension is zero the ETW brane cuts off half of the AdS space and when $T=L^{-1}$ the ETW brane coincides with the conformal boundary. The limit of $T\lesssim\frac{1}{L}$, i.e., the limit in which the brane approaches the “would be conformal boundary” is referred to as the critical limit. The critical limit is physically relevant in the discussion of gravity induced on the brane, particularly in higher dimensions. It is a limit in which bulk gravitons become localized on the brane and one can consider a conventional lower dimensional gravity theory on the world volume of the brane Karch:2000ct ; Geng:2020qvw ; Neuenfeld:2021wbl . In the three-dimensional case discussed here, a theory of dilaton gravity gets induced on the brane, and the critical limit is the limit in which the cutoff scale of the brane theory becomes small Chen:2020uac ; Geng:2022slq ; Geng:2022tfc ; Neuenfeld:2023svs . In the case of Dirichlet BCs we simply interpret the brane as some cutoff surface of constant extrinsic curvature $K=2L^{-1}\sin\theta_{0}$, in this perspective there is no dynamical gravity on the brane [^5] With Dirichlet BCs the brane is more similar to a cutoff surface discussed in the $T\bar{T}$ context, whereas Neumann BCs is more appropriate in the discussion of islands in double holography and dynamical gravity on the brane.. In either perspective the key parameter which dictates how the brane/cutoff is embedded in the bulk is given by $\theta_{0}$ and all our results will be formulated in terms of $\theta_{0}$.
@@ 2.3 General Classification of Extremal Surfaces in Poincare AdS3
Nonetheless, for $|k|$ above a certain threshold, the extremal surface becomes timelike. Here, we consider only spacelike separated points with respect to the bulk metric, as expected in computations of entanglement entropy for spacelike separated subregions. [^7] For work on timelike entanglement entropy see Doi:2022iyj ; Liu:2022ugc ; Doi:2023zaf ; Li:2022tsv ; Narayan:2022afv ; Narayan:2023zen . In this case we have that $|\Delta t|<\sqrt{\Delta x^{2}+\Delta z^{2}}$, where $\Delta x$ and $\Delta z$ are the coordinate distances between the RT endpoints. This places an upper bound on the parameter $|k|<k_{c}$, with $k_{c}$ given by
@@ 3.1 Condition for EWN
where the right hand side defines the additional bounds $\Delta t_{i,EWN}$ [^11] Among the two bounds the $i=1$ bound is more stringent. However, as we will soon show, satisfying $i=1$ bound is not necessary to satisfy EWN. Although the $i=1$ constraint is not relevant when formulating necessary and sufficient conditions for EWN it will reappear in our discussion of trying to understand non-locality in cutoff holography in Section 3.4..
@@ 3.4 Consequences for Locality in Cutoff Holography
In addition, it is also interesting to consider what the condition $|\Delta t|<\Delta t_{A,EWN}$ implies about the separation between $A$ and $\text{Vir}(B)$ (which was used in the discussion of Eq. (23)). Taking inspiration from recent works which think about a notion of induced causality on the brane in terms of subregions that exist behind the brane on the imaginary conformal boundary Mori:2023swn ; Franken:2024wmh , we will think of $B$ as a coarse grained version of $\text{Vir}(B)$ which lives on the imaginary conformal boundary behind the brane. Through this identification we might suggest a natural way to think about independent subregions on the brane/cutoff is to restrict $B$ to be the such that $\text{Vir}(B)$ is spacelike separated from $A$ in the full spacetime without the brane. Interestingly enough, this exactly gives the condition that $|\Delta t|<\Delta t_{1,EWN}$ which was derived as a sufficient condition for EWN in Eq. (36) [^19] Recall that in Section 3.1, this condition represented the most stringent sufficient condition to ensure $\mathcal{W}_{E}(B)\subseteq\mathcal{W}_{E}(A\cup B)$, we disregarded it because it was actually not a necessary condition but it reappears here as well.. Using the fact that $\Delta t_{2,EWN}-\Delta t_{1,EWN}=2a_{-}$ along with the identity in Eq. (47) we can easily see that that $\Delta t_{1,EWN}-\Delta t_{A,EWN}=-\ell_{*}\leq 0$, where saturation occurs when $\theta_{0}=\pi/2$. This implies that any time $A$ and $\text{Vir}(B)$ are spacelike separated then EWN is will be satisfied. However, due to the fact that $\Delta t_{1,EWN}-\Delta t_{A,EWN}\leq 0$ there also exist configurations where EWN is satisfied but $A$ and $\text{Vir}(B)$ fail to be spacelike separated and therefore would not be independent in this alternate notion of subregion independence formulated in terms of subregions behind the brane.
@@ 4 Entanglement Wedge Nesting in a Planar BTZ Black Hole Geometry
In this section, we will study conditions for entanglement wedge nesting for a two-sided planar BTZ black hole geometry with one exterior having an ETW brane by adapting the formalisms developed in the previous sections in AdS3. BTZ black hole geometry has also been of great interest in a variety of contexts ranging from black hole microstates Hartman:2013qma ; Kourkoulou:2017zaj ; Miyaji:2021ktr , brane world cosmologies Cooper:2018cmb ; Antonini:2019qkt ; Waddell:2022fbn , as well as constraining DGP couplings in brane-world setups Lee:2022efh ; Geng:2022slq ; Geng:2022tfc ; Geng:2023qwm , which motivates their study using our techniques.
@@ 4.1 Two-Sided BTZ Black Holes with an ETW Brane
We now introduce an ETW brane in our setup. The ETW brane will be a timelike co-dimension one surface, which satisfies the following equations of motion involving the extrinsic curvature and induced metric [^20] As we discussed in the introduction, this is when we impose Neumann BCs on the brane. In the Dirichlet case we simply view the brane as a cutoff with constant extrinsic curvature set by $y_{\text{Br}}$. All our results will be formulated in terms of $y_{\text{Br}}$.
@@ 4.4 Conditions for EWN
The tightest interval occurs when we take the limit of the expressions above as $\Xi\to 1$ (i.e. when $a\to\infty$) to give [^22] In Appendix C.6 we show that the connected phase will dominate in this regime. Intuitively, the thermal extremal surface area goes to infinity while the connected surface area remains finite as $a\to\infty$.,